OpenStudy (anonymous):

a certain sum is divided among a,b,c in such a way that a gets 3 times as much as b, b receives 2 times as much as c. what is the sum divided ?

OpenStudy (nowhereman):

So in formula you have \(s = a+b+c\) and \(b = 2c\), \(a = 3b = 6c\) so \[s = 6c + 2c + c = 9c\] so \[c = \frac{1}{9} s\]\[b = \frac{2}{9}s\]\[a = \frac{2}{3}s\]

OpenStudy (anonymous):

let c gets x so b gets 2x so a gets 6x now x+2x+3x=sum x=sum/9 so a gets 2*sum/3 b=2*sum/9 a=sum/9

OpenStudy (anonymous):

oppss !! i have not written full quest a certain sum is divided among a,b,c in such a way that a gets 3 times as much as b, b receives 2 times as much as c. IF A RECEIVES 1500 MORE THAN C. what is the sum divided ?

OpenStudy (anonymous):

sorry for incovienece to all !! :((

OpenStudy (anonymous):

ok.. a receives 2*sum/3, c receives sum/9 so 2sum/3-sum/9=1500 5sum/9=1500 sum=2700 a gets 1800, b=600, c=300

OpenStudy (anonymous):

thanks nowhereman...for medal..